# L(M)V-IQL: Multiple Intention Inverse Reinforcement Learning for Animal Behavior Characterization

### v1

@@ 4.1 Clustering of Independent Latent States
We start from the simpler case where the occurrence of different intentions satisfies a generalized Bernoulli process. Let $\{\nu_{1},\ldots,\nu_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ be the set of prior probability corresponding to the occurrence of each latent state $z_{k}$, the set of parameters $\Theta$ to be inferred is then $\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$. The optimal value for respective parameters in this parameter set at each EM iteration is provided by Theorem 1:
Theorem 1.
Given that the intention transition dynamics satisfies a generalized Bernoulli process, at iteration $\tau$, the EM update equation (Equation 4) for each parameter in the corresponding parameter set $\Theta=\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ is given by
$&\nu_{k}^{\tau+1}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}\end{aligned}\right.$ (5)
for all $\nu,{\mathcal{R}}\in\Theta$, where $\zeta_{ik}\coloneqq\frac{1}{Z}\nu_{k}\prod_{(s,a)\in\xi_{i}}\pi_{{\mathcal{R}}_{k}}(s,a)$ is the probability that trajectory $i$ was demonstrated under latent state $z_{k}$ normalized by factor $Z$.
@@
See Appendix A.1. ∎
Noting that updating the reward function estimation ${\mathcal{R}}$ at each iteration according to Equation 5 is equivalent to solving Problem 1, except that each trajectory is weighted by a probability $\zeta$ during sampling. Thus combining the above EM approach for trajectory clustering with IAVI or IQL algorithms leads to the class of Latent Variable Inverse Q-learning (LV-IQL) algorithms (cf. Algorithm 1), solving the MI-IQL problem (Problem 2) when the occurrence of different latent states is independent.
@@ 4.2 Clustering of Markovian Interdependent Latent States
In addition to the generalized Bernoulli process, the Markov process is also considered an alternative for describing intention transition dynamics (Ashwood et al., 2022b; Le et al., 2023), where the occurrence of the next latent state corresponding to a specific intention is dependent and only dependent on the previous latent state. Under this assumption, given the agent demonstrations ${\mathcal{D}}$ consisting of a sequence of trajectories, the set of parameters to be inferred is then $\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$, where $\Pi\colon\Delta({\mathcal{Z}})$ and $\Lambda\colon{\mathcal{Z}}\to\Delta({\mathcal{Z}})$ ($\Delta$ is the probability simplex) denoting the latent state initial distribution probability and latent state transition matrix, respectively. The optimal value for respective parameters in the parameter set $\Theta$ at each EM iteration is provided by Theorem 2:
Theorem 2.
Given that the intention transition dynamics satisfies a Markov process, at iteration $\tau$, the EM update equation (Equation 4) for each parameter in the corresponding parameter set $\Theta=\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ is given by [^1] Here we assume the index of trajectories in ${\mathcal{D}}$ starts from 0 instead of 1 for convenience but without losing generality.
$&\pi_{k}^{\tau+1}\coloneqq\Pr(y_{0}=k\mid{\mathcal{D}},\Theta^{\tau})\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid{\mathcal{D}},\Theta^{\tau})}{\sum_{i=1}^{N}\Pr(y_{i-1}=k\mid{\mathcal{D}},\Theta^{\tau})}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=0}^{N}\Pr(y_{i}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}\end{aligned}\right.$ (6)
for all $\pi,\Lambda,{\mathcal{R}}\in\Theta$.
@@
See Appendix A.2. ∎
In practice, the Forward-Backward algorithm (Baum et al., 1970) can be used to address the probabilities in Equation 6, and IAVI or IQL then estimates the corresponding reward function independently for each latent state. This leads to the class of Latent Markov Variable Inverse Q-learning (LMV-IQL) algorithms (cf. Algorithm 2, details see also Appendix B.2), which solves Problem 2 when the latent state transition satisfies the Markov property.
@@ 5.1 Application of LV-IQL to Simulated Behavior
Figure 1: Gridworld environment architecture with agent trajectories.
@@
We compared between the performance of LV-IAVI, LV-IQL, IAVI, and IQL trained on the whole demonstration space. Two latent states were considered for both LV-IAVI and LV-IQL. As a measure of performance, we used the Expected Value Difference (EVD) metric (Levine et al., 2011). EVD is defined as the mean square error between the state-value under the true reward function for the expert policy and the state-value under the true reward for the optimal Boltzmann policy w.r.t. the learnt reward. It provides an estimation of the sub-optimality of the learnt policy under the true reward function. For LV-IAVI and LV-IQL, the inferred latent states with respective trajectory clusters were assigned to the best-fit ground truth intentions. Since IAVI and IQL assumed all trajectories were demonstrated under one intention, the EVD was analyzed twice on the ground truth reward for different intentions with the same learnt Boltzmann policy (Figure 2). The trajectory clusters learnt with LV-IAVI and LV-IQL are highly overlapped with the ground truth trajectory distribution. As a result, the learnt reward functions via LV-IAVI and LV-IQL match the respective ground truth reward functions exactly, while the single intention IAVI and IQL only resulted in a large EVD of $\sim 21$ for the ‘hungry’ intention and $\sim 38$ for the ‘thirsty’ intention, representing a mixed reward function for the two intentions. Similar results were found when we removed the punishment on intention irrelevant rewards (Appendix C.2). Noting that the DIRL algorithm (Ashwood et al., 2022a) is infeasible here as it assumes continuously time-varying rewards, which only addresses cases where the intention transition occurs after each action execution. However, in the above Gridworld experiment, each episode was conducted under one of two intentions, where the intention remains fixed within the episode, and the transition between intentions only occurs between episodes.
@@ 5.2 Application of LMV-IAVI to Mice Navigating Trajectories
We demonstrate model comparison by first applying our method to trajectories from the water-restricted animals. The test set log-likelihood (LL) is similar for LMV-IAVI and DIRL under single intention ($K=1$). However, LMV-IAVI with $K>2$ substantially outperforms DIRL (Figure 3). Although the test LL continues to grow for larger $K$, the Bayesian information criterion (BIC) appears to increase (Figure 3). Thus LMV-IAVI with 2 latent states is considered for subsequent analysis. The learnt mice policy under latent state 1 (‘Tired’) displays a preference of moving out from the water port towards the maze entrance and stay, while the policy under latent state 2 (‘Thirsty’) guides the mice directly to the water port along the optimal track. Correspondingly, in the ‘Tired’ latent state, the highest state occupancy is noted at the entrance state, while under ‘Thirsty’, it is observed at the water port (Figure 3). To delve into the intention transition dynamics, we computed the posterior probability over mice’s latent state across all trajectories. The recovered average temporal latent state trajectories show a high probability of the ‘Thirsty’ latent state at the beginning but later on tailed off, as the ‘Tired’ latent state gradually became dominant (Figure 3). These findings demonstrate that LMV-IAVI not only excels the state-of-the-art in predicting mice labyrinth navigating behavior, but also provides distinct and interpretable reward functions. Similarly, our LMV-IAVI algorithm again outperforms DIRL when applied to the water-unrestricted animal trajectory dataset. Further details can be found at Appendix D.2.
@@ 5.3 Application of LMV-IQL to Mice Reversal-learning Behavior
The inferred mice policies from LMV-IQL define how the subjects make decisions under three intentions (Figure 4). One of these policies, operating within latent state 1, displays a strong inclination toward adopting a ‘win-stay’ and ‘lose-switch’ strategy, which is the optimal policy in this deterministic reward bandit task. On the other hand, within latent state 2, the policy, referred to as the ‘Win-stay’ policy, exhibits a preference for exploitation when the previous trial was successful. However, following error trials, it employs a random action selection strategy, indicated by a $\sim 0.5$ probability of executing a ‘lose-switch’. Lastly, in latent state 3, a characteristic ‘Exploration’ policy emerges, where the subject consistently favors selecting the option opposite to the one chosen in the preceding trial, irrespective of whether they had won or lost in that particular instance. The recovered latent state trajectories in the example session reveals that the most probable latent state often exhibits a probability close to 1, indicating a high degree of confidence in discerning the subject’s intent based on the observed data (Figure 5). The ‘Exploration’ intention predominantly manifested at the onset of a block and endured for a relatively brief duration, in alignment with the learned latent state transition matrix (Figure 4).
@@
The significant values along the diagonal of the transition matrix within latent states 1 and 2, corresponding to ‘Exploitation’ and ‘Win-stay’, signify a heightened preference for persisting in the same latent state over multiple consecutive trials. Additionally, it becomes evident that error trials tend to coincide with the trials where the posterior probability of ‘Win-stay’ and ‘Exploration’ latent states reaches its zenith, corroborating the presence of suboptimal exploratory behavior associated with these two intentions. The average latent state transition trajectories across all blocks closely resembles those observed in the example session (Figure 4). As each block begin with the animal’s performance at a relatively low level, there is a decline in the posterior probability associated with the ‘Exploitation’ latent state, accompanied by an increase in the probabilities of the other two latent states associated with suboptimal exploratory strategies. Nonetheless, as the subjects’ performance steadily improves, the ‘Exploitation’ latent state progressively reasserts its dominance. Finally, to quantify latent state occupancies across all sessions, we assigned each trial to its most probable state. In contrast to the cohort’s general correct rate of $0.74\pm 0.02$, mice performed significantly better within the ‘Exploitation’ latent state, achieving a correctness rate of $0.86\pm 0.01$. In comparison, they attained lower correctness rates of $0.64\pm 0.03$ and $0.56\pm 0.06$ in the two alternative latent states (Figure 4). Furthermore, it’s worth noting that the mean posterior probability of the ‘Exploration’ latent state at the beginning of a new block shows a positive correlation with the average probability of the ‘Exploitation’ state at the end of the preceding block (Figure 4), suggesting that the ‘Exploration’ latent state appears to involve a deliberate, exploration-oriented action selection when mice are highly engaged and possess a good understanding of the environment.
@@ 6 Conclusion
In this study, we introduce a novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms for characterizing animal behavior during complex decision-making tasks. We extend the class of IQL algorithms (Kalweit et al., 2020) to learn multiple discrete reward functions from demonstrations. Specifically, we address the two most prevalent types of intention transition dynamics: the generalized Bernoulli process and the Markov process, under both model-based and model-free contexts. To validate our framework and compare with the state-of-the-art, we conduct experiments on simulated and real animal behavior data. Our approaches demonstrate a substantial improvement in behavior prediction compared to DIRL (Ashwood et al., 2022a) on mice navigation trajectories (indicated by the LL on held-out trajectories), without losing interpretability of the learnt reward functions. Moreover, our method provides distinct and interpretable reward functions for the mice cohort engaged in the reversal-learning task, where the animals displayed a pattern of alternating between exploitation and exploration intentions, which could extend over several consecutive trials within a single session. The transitions between these intentions followed a typical block-correlated trajectory, wherein the mice were more likely to exhibit in exploratory behaviors at the start of a new block, particularly if they had been highly engaged in the task in the preceding block.
A compelling avenue for future research lies in extending our framework to involve function approximations, which would enable the learning of a low-dimensional embedding of each state in the environment via e.g. a deep neural network. Such extension would allow us to scale our approach to high-dimensional or continuous state spaces, while also enabling the generalization across states. Another promising direction would be to extend the fixed intention transition probabilities with e.g. a generalized linear model, to incorporate the identification of potential external factors that influence intention transition dynamics.
Acknowledgments

### v2

@@ 4.1 Clustering of Independent Latent States
We start from the simpler case where the occurrence of different intentions satisfies a generalized Bernoulli process. Let $\{\nu_{1},\ldots,\nu_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ be the set of prior probability corresponding to the occurrence of each latent state $z_{k}$, the set of parameters $\Theta$ to be inferred is then $\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$. The optimal value for respective parameters in this parameter set at each EM iteration is provided by Theorem 4.2.
Theorem 4.2.
Given that the intention transition dynamics satisfies a generalized Bernoulli process, at iteration $\tau$, the EM update equation (Equation 5) for each parameter in the corresponding parameter set $\Theta=\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ is given by
$&\nu_{k}^{\tau+1}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\\ &r_{k}^{\tau+1}\coloneqq\operatorname*{argmax}_{r_{k}}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}\end{aligned}\right.$ (6)
for all $\nu,r\in\Theta$, where
$\zeta_{ik}\coloneqq\frac{1}{Z}\nu_{k}\prod_{(s,a)\in\xi_{i}}\pi_{r_{k}}(s,a)$ (7)
is the probability that trajectory $i$ was demonstrated under latent state $z_{k}$ normalized by factor $Z$.
@@
See Section A.1. ∎
Noting that updating the reward function estimation $r$ at each iteration according to Equation 6 is equivalent to solving 3.1, except that each trajectory is weighted by a probability $\zeta$ during sampling. Thus combining the above EM approach for trajectory clustering with IAVI or IQL algorithms leads to the class of Latent Variable Inverse Q-learning (LV-IQL) algorithms (cf. Algorithm 1), solving the MI-IQL problem (4.1) when the occurrence of different latent states is independent.
@@ 4.2 Clustering of Markovian Interdependent Latent States
In addition to the generalized Bernoulli process, the Markov process is also considered an alternative for describing intention transition dynamics (Ashwood et al., 2022b; Le et al., 2023). Under this assumption, given the agent demonstrations $\mathcal{D}$ consisting of a sequence of trajectories, the set of parameters to be inferred is then $\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$, where $\Pi\colon\mathcal{P}(\mathcal{Z})$ and $\Lambda\colon\mathcal{Z}\to\mathcal{P}(\mathcal{Z})$ ($\mathcal{P}$ is the probability simplex) denoting the latent state initial distribution probability and latent state transition matrix, respectively. The optimal value for respective parameters in the parameter set $\Theta$ at each EM iteration is provided by Theorem 4.3.
Theorem 4.3.
Given that the intention transition dynamics satisfies a Markov process, at iteration $\tau$, the EM update equation (Equation 5) for each parameter in the corresponding parameter set $\Theta=\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$ is given by
$&\Pi_{k}^{\tau+1}\coloneqq\Pr(y_{0}=k\mid\mathcal{D},\Theta^{\tau})\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid\mathcal{D},\Theta^{\tau})}{\sum_{i=1}^{N}\Pr(y_{i-1}=k\mid\mathcal{D},\Theta^{\tau})}\\ &r_{k}^{\tau+1}\coloneqq\operatorname*{argmax}_{r_{k}}\sum_{i=0}^{N}\Pr(y_{i}=k\mid\mathcal{D},\Theta^{\tau})\\ &\qquad\qquad\times\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}\end{aligned}\right.$ (8)
for all $\Pi,\Lambda,r\in\Theta$ [^1] Here we assume the index of trajectories in $\mathcal{D}$ starts from 0 instead of 1 for convenience without losing generality..
@@
See Section A.2. ∎
In practice, the Forward-Backward algorithm (Baum et al., 1970) can be used to address the probabilities in Equation 8, and IAVI or IQL then estimates the corresponding reward function independently for each latent state. This leads to the class of Latent Markov Variable Inverse Q-learning (LMV-IQL) algorithms (cf. Algorithm 2, details see also Section B.2), which solves 4.1 when the latent state transition satisfies the Markov property.
@@ 5.1 Application of LV-IQL to Simulated Behavior
Figure 1: Gridworld environment with demonstrated trajectories.
@@
We compared between the performance of LV-IAVI, LV-IQL, IAVI, and IQL trained on the whole demonstration space. Two latent states were considered for both LV-IAVI and LV-IQL. As a measure of performance, we used the Expected Value Difference (EVD) metric (Levine et al., 2011). EVD is defined as the mean square error between the state-value under the true reward function for the expert policy and the state-value under the true reward for the optimal Boltzmann policy w.r.t. the learnt reward. It provides an estimation of the sub-optimality of the learnt policy under the true reward function. For LV-IAVI and LV-IQL, the inferred latent states with respective trajectory clusters were assigned to the best-fit ground truth intentions. Since IAVI and IQL assumed all trajectories were demonstrated under one intention, the EVD was analyzed twice on the ground truth reward for different intentions with the same learnt Boltzmann policy (Figure 2). The trajectory clusters learnt with LV-IAVI and LV-IQL are highly overlapped with the ground truth trajectory distribution. As a result, the learnt reward functions via LV-IAVI and LV-IQL match the respective ground truth reward functions exactly, while the single intention IAVI and IQL only resulted in a large EVD of $\sim 21$ for the ‘hungry’ intention and $\sim 38$ for the ‘thirsty’ intention, representing a mixed reward function for the two intentions. Similar results were found when we removed the punishment on intention irrelevant rewards (Section C.2). Noting that the DIRL algorithm (Ashwood et al., 2022a) is infeasible here as it assumes continuously time-varying rewards, which only addresses cases where the intention transition occurs after each action execution. However, in the above Gridworld experiment, each episode was conducted under one of two intentions, where the intention remains fixed within the episode, and the transition between intentions only occurs between episodes.
@@ 5.2 Application of LMV-IAVI to Mice Navigating Trajectories
We demonstrate model comparison by first applying our method to trajectories from the water-restricted animals. The test set log-likelihood (LL) is similar for LMV-IAVI and DIRL under single intention ($K=1$). However, LMV-IAVI with $K>2$ substantially outperforms DIRL (Figure 3). Although the test LL continues to grow for larger $K$, the Bayesian information criterion (BIC) appears to increase (Figure 3). Thus LMV-IAVI with 2 latent states is considered for subsequent analysis. The learnt mice policy under latent state 1 (‘Tired’) displays a preference of moving out from the water port towards the maze entrance and stay, while the policy under latent state 2 (‘Thirsty’) guides the mice directly to the water port along the optimal track. Correspondingly, in the ‘Tired’ latent state, the highest state occupancy is noted at the entrance state, while under ‘Thirsty’, it is observed at the water port (Figure 3). To delve into the intention transition dynamics, we computed the posterior probability over mice’s latent state across all trajectories. The recovered average temporal latent state trajectories show a high probability of the ‘Thirsty’ latent state at the beginning but later on tailed off, as the ‘Tired’ latent state gradually became dominant (Figure 3). These findings demonstrate that LMV-IAVI not only excels the state-of-the-art in predicting mice labyrinth navigating behavior, but also provides distinct and interpretable reward functions. Similarly, our LMV-IAVI algorithm again outperforms DIRL when applied to the water-unrestricted animal trajectory dataset. Further details can be found at Section D.2.
@@ 5.3 Application of LMV-IQL to Mice Reversal-learning Behavior
The inferred mice policies from LMV-IQL define how the subjects make decisions under three intentions (Figure 4). One of these policies, operating within latent state 1, displays a strong inclination toward adopting a ‘win-stay’ and ‘lose-switch’ strategy, which is the optimal policy in this deterministic reward bandit task. On the other hand, within latent state 2, the policy, referred to as the ‘Win-stay’ policy, exhibits a preference for exploitation when the previous trial was successful.
@@
However, following error trials, it employs a random action selection strategy, indicated by a $\sim 0.5$ probability of executing a ‘lose-switch’. Lastly, in latent state 3, a characteristic ‘Exploration’ policy emerges, where the subject consistently favors selecting the option opposite to the one chosen in the preceding trial, irrespective of whether they had won or lost in that particular instance. The recovered latent state trajectories in the example session reveals that the most probable latent state often exhibits a probability close to 1, indicating a high degree of confidence in discerning the subject’s intent based on the observed data (Figure 5). The ‘Exploration’ intention predominantly manifested at the onset of a block and endured for a relatively brief duration, in alignment with the learned latent state transition matrix (Figure 4). The significant values along the diagonal of the transition matrix within latent states 1 and 2, corresponding to ‘Exploitation’ and ‘Win-stay’, signify a heightened preference for persisting in the same latent state over multiple consecutive trials. Additionally, it becomes evident that error trials tend to coincide with the trials where the posterior probability of ‘Win-stay’ and ‘Exploration’ latent states reaches its zenith, corroborating the presence of suboptimal exploratory behavior associated with these two intentions. The average latent state transition trajectories across all blocks closely resembles those observed in the example session (Figure 4). As each block begin with the animal’s performance at a relatively low level, there is a decline in the posterior probability associated with the ‘Exploitation’ latent state, accompanied by an increase in the probabilities of the other two latent states associated with suboptimal exploratory strategies. Nonetheless, as the subjects’ performance steadily improves, the ‘Exploitation’ latent state progressively reasserts its dominance. Finally, to quantify latent state occupancies across all sessions, we assigned each trial to its most probable state. In contrast to the cohort’s general correct rate of $0.74\pm 0.02$, mice performed significantly better within the ‘Exploitation’ latent state, achieving a correctness rate of $0.86\pm 0.01$. In comparison, they attained lower correctness rates of $0.64\pm 0.03$ and $0.56\pm 0.06$ in the two alternative latent states (Figure 4). Furthermore, it’s worth noting that the mean posterior probability of the ‘Exploration’ latent state at the beginning of a new block shows a positive correlation with the average probability of the ‘Exploitation’ state at the end of the preceding block (Figure 4), suggesting that the ‘Exploration’ latent state appears to involve a deliberate, exploration-oriented action selection when mice are highly engaged and possess a good understanding of the environment.
@@ 6 Conclusion
In this study, we introduce a novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms for characterizing animal behavior during complex decision-making tasks. We extend the class of IQL algorithms (Kalweit et al., 2020) to learn multiple discrete reward functions from demonstrations. Specifically, we address the two most prevalent types of intention transition dynamics: the generalized Bernoulli process and the Markov process, under both model-based and model-free contexts.
To validate our framework and compare with the state-of-the-art, we conduct experiments on simulated and real animal behavior data. Our approaches demonstrate a substantial improvement in behavior prediction compared to DIRL (Ashwood et al., 2022a) on mice navigation trajectories (indicated by the LL on held-out trajectories), without losing interpretability of the learnt reward functions. Moreover, our method provides distinct and interpretable reward functions for the mice cohort engaged in the reversal-learning task, where the animals displayed a pattern of alternating between exploitation and exploration intentions, which could extend over several consecutive trials within a single session. The transitions between these intentions followed a typical block-correlated trajectory, wherein the mice were more likely to exhibit in exploratory behaviors at the start of a new block, particularly if they had been highly engaged in the task in the preceding block.
A compelling avenue for future research lies in extending our framework to involve function approximations, which would enable the learning of a low-dimensional embedding of each state in the environment via e.g. a deep neural network. Such extension would allow us to scale our approach to high-dimensional or continuous state spaces, while also enabling the generalization across states. Another promising direction would be to extend the fixed intention transition probabilities with e.g. a generalized linear model (Nguyen et al., 2015), to incorporate the identification of potential external factors that influence intention transition dynamics. Finally, we also plan to extend our L(M)V-IQL algorithm to accommodate an unknown number of reward functions (Michini & How, 2012; Surana & Srivastava, 2014).
Broader Impact
Our work constitutes an advancement in IRL methods that can be used in behavioral neuroscience and cognitive science researches. Importantly, our work can be adapted to multiple behavioral experiment paradigms with minor modifications and enables the implementation by non-expert users. This is expected to boost productivity and tighten interdisciplinary research collaborations. On the other hand, we believe that our work also provides future directions in machine learning and robotics, such as the development of advanced reinforcement learning algorithms. By gaining insights into the intrinsic reward function driving animal behavior, we anticipate that our framework could serve as a valuable resource for formulating improved reward functions when training artificial agents or robots to perform challenging tasks. This can potentially accelerate existing trends for automation in industries.
Acknowledgements